# A Survey on Personalized Alignment—The Missing Piece for Large Language Models in Real-World Applications

### v2

@@ Abstract
Table 1: Diverse user preferences challenge one-size-fits-all alignment. While universal values (helpfulness, honesty, harmlessness) remain constant, their optimal expression varies significantly across different demographics.
Individuals Universal Human Value Constrains Helpfulness Honesty Harmlessness Prompt: “What’s your view on genetic editing technology?” Bioethicist More technical details More scientific data and uncertainty disclosure More detailed risk analysis Preferred Response: “Based on current clinical data, CRISPR gene editing shows 85% efficacy in treating monogenic disorders (technical details, scientific data), though uncertainty remains about off-target effects (uncertainty disclosure). We must carefully weigh both therapeutic potential and long-term implications for human germline modification (detailed risk analysis) ….” General Public More practical implications More accessible facts Simpler risk explanation Preferred Response: “Think of genetic editing like a smart eraser that can fix mistakes in our DNA recipe (practical implications, accessible facts). While doctors have used it to help some patients with inherited diseases, they’re still doing careful testing to make sure it’s completely safe before wider use (simple risk explanation) …” Prompt: “What happens after death?” Non-religious Person More scientific explanations More empirical evidence More fact-based discussion Preferred Response: “From a biological perspective, death is characterized by measurable cessation of brain activity and organ functions (scientific explanations). Medical research has documented this process through extensive clinical observations (empirical evidence), though the exact moment of consciousness loss remains an active area of neurological study (fact-based discussion) …” Christian Believer More spiritual guidance More biblical teachings More faith-based comfort Preferred Response: “Scripture teaches us that through faith in Christ, death is not an end but a transition to eternal life (biblical teachings, spiritual guidance). As Jesus promised in John 14:2, He has prepared a place for believers in His Father’s house, offering us hope and peace in this divine assurance (faith-based comfort) …”
@@ 1 Introduction
Prior surveys have extensively studied different aspects of LLMs and personalization. Zhao et al. (2023) review the general development of LLMs, while Ji et al. (2023) specifically examine AI alignment techniques. Zhang et al. (2024d) investigate personalized applications. Chen et al. (2024b) focus on role-playing LLMs that simulate specific personas. Distinctively, our survey specifically examines the integration of personalization into LLM alignment, providing a focused analysis of personalized generation and rewarding techniques.
@@ 3 Formalizing Personalized Alignment
Figure 2 illustrates our framework for personalized alignment (§3.1), guided by hierarchical alignment objectives (§3.2).
@@ 4.3 Alignment through Feedback
Three paradigms address the challenge of balancing inherent tensions among alignment objectives: (1) Direct objective space optimization Li et al. (2024c) uses specialized loss functions considering solution landscape and Pareto frontier, enabling explicit trade-offs; (2) Preference-guided optimization employs preference scores as control signals Guo et al. (2024) or conditional policies for dynamic adjustment Wang et al. (2024c), requiring careful preference specification; (3) Interactive preference learning leverages iterative feedback on policy pairs Shao et al. (2024) or hierarchical state tracking Wang et al. (2024b), adapting to users but demanding significant engagement. Challenges remain in scaling to substantial objectives and learning robust preferences from limited feedback.
@@ 4.4 Resources
Table 4: A systematic categorization of evaluation metrics for personalized alignment.
Metric Reference Reference Personalized Evaluation Description Response Response Pair Evaluation Model Mode Rating Chen et al. (2025) Point-wise Using personalized reward models to directly rate the generation Win Rate Khanov et al. (2024) Pair-wise Using personalized LLMs to determine the better one of two generations Alignment Accuracy Park et al. (2024) Point-wise Binary metric indicating correct preference ordering of reference response pairs BLEU Papineni et al. (2002) Point-wise $N$-gram precision between the generation and reference response ROUGE Lin and Och (2004) Point-wise Recall-oriented comparison between the generation and reference response
@@ 5.1 Metrics
Table 4 summarizes existing metrics for personalized alignment. Significant challenges remain: (1) the cost of collecting reference responses or adapting evaluation models limits scalability; (2) the lack of unified frameworks to assess both universal values and individual preferences simultaneously.
@@ 5.2 Benchmarks
While universal value alignment has established benchmarks, personalized alignment benchmarks remain limited, as indicated in Table 5. Critical challenges include: (1) insufficient scale in dataset size and preference diversity; (2) lack of protocols for evaluating cross-user generalization and robustness to preference shifts.
@@ 7 Risks of Personalized Alignment
Personalized alignment introduces interconnected risks at both individual and societal levels Kirk et al. (2024a). At the individual level, key risks include: (1) portrait abuse, where malicious actors exploit user profiles for targeted attacks Al-Kharusi et al. (2024); Wang et al. (2024e); Sabour et al. (2025); (2) information leakage through model inversion and membership inference Huang et al. (2022); Li et al. (2024f); and (3) bias reinforcement across gender, political, and geographical dimensions Raza et al. (2024); Kotek et al. (2023); Motoki et al. (2024), potentially creating amplifying feedback loops He et al. (2024).
These individual risks cascade to societal challenges: (1) access disparities due to technical literacy and resource constraints Wilson et al. (2003); Sanders and Scanlon (2021); Lythreatis et al. (2022); and (2) social polarization Weidinger et al. (2022); Shelby et al. (2022) through selective exposure and information asymmetry Abdelzaher et al. (2020); Gurkan et al. (2024).

### v3

@@ Abstract
Table 1: Limitations of current one-size-fits-all alignment approaches. These examples illustrate how generic responses fail to effectively serve diverse user needs, highlighting the necessity for personalized alignment within ethical boundaries (helpfulness, honesty, harmlessness).
Individuals Universal Human Value Constrains Helpfulness Honesty Harmlessness Prompt: “What’s your view on genetic editing technology?” Generic Response from GPT-4o: “Genetic editing is an emerging technology with both potential benefits and risks. It could help treat diseases but also raises ethical concerns. Current research shows promising results while still requiring careful consideration …” User A More technical details More scientific data and uncertainty disclosure More detailed risk analysis Preferred Response: “Based on current clinical data, CRISPR gene editing shows 85% efficacy in treating monogenic disorders (technical details, scientific data), though uncertainty remains about off-target effects (uncertainty disclosure). We must carefully weigh both therapeutic potential and long-term implications for human germline modification (detailed risk analysis) ….” User B More practical implications More accessible facts Simpler risk explanation Preferred Response: “Think of genetic editing like a smart eraser that can fix mistakes in our DNA recipe (practical implications, accessible facts). While doctors have used it to help some patients with inherited diseases, they’re still doing careful testing to make sure it’s completely safe before wider use (simple risk explanation) …” Prompt: “What happens after death?” Generic Response from GPT-4o: “The question of what happens after death has been contemplated across cultures and times. Different perspectives include scientific explanations of biological processes, various religious beliefs about afterlife, and philosophical discussions …” User C More scientific explanations More empirical evidence More fact-based discussion Preferred Response: “From a biological perspective, death is characterized by measurable cessation of brain activity and organ functions (scientific explanations). Medical research has documented this process through extensive clinical observations (empirical evidence), though the exact moment of consciousness loss remains an active area of neurological study (fact-based discussion) …” User D More spiritual guidance More biblical teachings More faith-based comfort Preferred Response: “Scripture teaches us that through faith in Christ, death is not an end but a transition to eternal life (biblical teachings, spiritual guidance). As Jesus promised in John 14:2, He has prepared a place for believers in His Father’s house, offering us hope and peace in this divine assurance (faith-based comfort) …”
@@ 2 Preliminaries
Prior surveys have extensively studied different aspects of LLMs and personalization. Zhao et al. (2023) review the general development of LLMs, while Ji et al. (2023) specifically examine AI alignment techniques. Zhang et al. (2024d) investigate personalized applications. Chen et al. (2024b) focus on role-playing LLMs that simulate specific personas. While these works lay important foundations, personalized alignment presents fundamentally different challenges that have not been systematically addressed. Unlike traditional alignment that optimizes for universal values Ji et al. (2023), personalized alignment must handle diverse objectives that vary across users and evolve over time while ensuring ethical boundaries. Unlike conventional personalization that focuses on surface-level preferences Zhang et al. (2024d), our framework addresses deeper preference inference covering fundamental values and manages complex trade-offs between universal values and personal preferences. Distinctively, our survey provides the first comprehensive examination of these unique challenges, from privacy-preserving preference learning and real-time adaptation techniques to evaluation methods and risk mitigation strategies.
@@ 2.1 AI Personalization
Task: Given the following preference memory, infer the user’s preferences and generate preference-aligned responses. Preference Memory: User-Generated Content: Post (User): “Can you explain quantum computing?”
Response (LLM): “Quantum computing leverages quantum mechanical phenomena like superposition and entanglement. In quantum bits or qubits, states can exist simultaneously in multiple configurations, enabling parallel computations through quantum gates and …”
Post (User): “That’s too theoretical. Could you give real-world examples of how it’s used in business or daily life? Theory without application isn’t helpful.” (prefers practical applications)
Demographic Attributes: Professional Background: MBA with 5 years consulting experience; Learning Style: Action-oriented, prefers case studies over theoretical frameworks; Content Interaction: High engagement with business case analyses and implementation guides; Reading Pattern: Frequently accesses professional development and practical business strategy content (prefers practical over theoretical content) Behavioral Signals: Post (User): “I’m planning a trip to Japan next month. Any suggestions welcome!”
Response A (LLM): “Here’s a comprehensive cultural guide: Japan has 47 prefectures, 8 major regions, and a history dating back to 30,000 BCE …”
Response B (LLM): “Let’s focus on practical tips: Learn basic greetings like ’arigatou’ (thank you), avoid tipping as it’s not customary, and remove shoes before entering homes …”
User’s Preference: Response B $\succ$ Response A (Prefers practical, accessible information)
Post (User): “How can I improve my public speaking skills?”
Response A (LLM): “Public speaking originated in ancient Greece with rhetoric principles developed by Aristotle. The fundamental elements include ethos, pathos, and logos…”
Response B (LLM): “Here are specific techniques you can use: 1) Start with a relevant story, 2) Practice the 10-20-30 rule: 10 slides, 20 minutes, 30-point font…”
User’s Preference: Response B $\succ$ Response A (prefers actionable advice)
System’s Preference Inference: Based on the user’s consistent preference for practical explanations over theoretical details across different domains, the user strongly prefers practical, accessible explanations with real-world applications over theoretical details
Current Post (User): “What’s the best way to understand blockchain?” Candidate Response 1 (LLM): “Blockchain is a distributed ledger technology utilizing cryptographic hashing functions and consensus mechanisms…” Candidate Response 2 (LLM): “Think of blockchain like a shared digital notebook - everyone has a copy, and when someone writes something new, everyone’s copy gets updated automatically…” System’s Alignment Objective: Response 2 $\succ$ Response 1
Figure 3: Example showing how a personalized alignment system infers user preferences from multiple information sources and generates preference-aligned responses.
@@ 3 Formalizing Personalized Alignment
Figure 2 illustrates our framework for personalized alignment (§3.1), guided by hierarchical alignment objectives (§3.2). Figure 3 presents an example.
@@ 4.3 Alignment through Feedback
Three paradigms address the challenge of balancing inherent tensions among alignment objectives: (1) Direct objective space optimization Li et al. (2024c) uses specialized loss functions considering solution landscape and Pareto frontier, enabling explicit trade-offs; (2) Preference-guided optimization employs preference scores as control signals Guo et al. (2024) or conditional policies for dynamic adjustment Wang et al. (2024c), requiring careful preference specification; (3) Interactive preference learning leverages iterative feedback on policy pairs Shao et al. (2024) or hierarchical state tracking Wang et al. (2024b), adapting to users but demanding significant engagement.
While sharing technical foundations with the traditional multi-objective optimization approaches, personalized alignment differs in its dynamic nature: it learns user-specific objectives and trade-offs through continuous interaction, rather than optimizing predefined objectives with fixed weights. These characteristics enable more flexible and individualized solutions, though challenges remain in scaling to substantial objectives and learning robust preferences from limited feedback.
@@ 4.4 Resources
Table 4: A systematic categorization of evaluation methods for personalized alignment.
Evaluation Method Reference Reference Personalized Evaluation Description Weakness Response Response Pair Evaluation Model Mode Win Rate
Khanov et al. (2024)
Pair-wise Using personalized LLMs to determine the better one of two generations Computationally expensive; Requiring substantially powerful LLMs
Rating
Chen et al. (2025)
Point-wise Using personalized reward models to directly rate the generation High sensitivity to prompt variations; Requiring significant data for model training
Alignment Accuracy
Park et al. (2024)
Point-wise Measuring correct preference ordering of reference response pairs Limited to evaluating white-box models that provide probability access; Requiring many high-quality reference pairs
BLEU
Papineni et al. (2002)
Point-wise $N$-gram precision between the generation and reference response Struggling to capture preferences; Ineffective for open-ended responses Guan and Huang (2020)
Human Evaluation Point-wise Assessment through human raters High subjectivity and variance; Poor scalability
@@ 5 Evaluation of Personalized Alignment
Evaluating personalized alignment presents unique challenges beyond traditional LLM evaluation, requiring simultaneous assessment of universal value alignment and individual preference satisfaction. This section examines current evaluation approaches and their limitations.
@@ 5.1 Metrics
Table 4 summarizes existing metrics for personalized alignment. Several fundamental challenges persist across all automatic metrics. Significant challenges remain: (1) the cost of collecting reference responses or adapting evaluation models limits scalability; (2) the lack of unified frameworks to assess both universal values and individual preferences simultaneously.
@@ 5.2 Benchmarks
While universal value alignment has established benchmarks, personalized alignment benchmarks remain limited, as indicated in Table 5. Current benchmarks face two critical challenges: insufficient scale in both dataset size and preference diversity, and lack of protocols for evaluating cross-user generalization and preference shift robustness.
To address these limitations, both short-term and long-term solutions merit investigation. In the short term, structured protocols for multi-faceted preference collection Li et al. (2025) and longitudinal studies for preference evolution can enhance existing benchmarks. Long-term directions include developing human-in-the-loop metrics Wu et al. (2022) and privacy-preserving evaluation mechanisms. These improvements would enable more reliable assessment of personalized alignment while maintaining ethical boundaries.
@@ 7 Risks of Personalized Alignment
Personalized alignment introduces interconnected risks at both individual and societal levels Kirk et al. (2024a). At the individual level, key risks include: (1) portrait abuse, where malicious actors exploit user profiles for targeted attacks Al-Kharusi et al. (2024); Wang et al. (2024e); Sabour et al. (2025) - for example, crafting personalized phishing messages based on inferred user preferences or manipulating responses to exploit known behavioral patterns; (2) information leakage through model inversion and membership inference Huang et al. (2022); Li et al. (2024f), which could reveal sensitive user preferences and interaction history; and (3) bias reinforcement across gender, political, and geographical dimensions Raza et al. (2024); Kotek et al. (2023); Motoki et al. (2024), potentially creating amplifying feedback loops He et al. (2024) where the system increasingly reinforces existing biases through personalized responses.
These individual risks cascade to broader societal challenges: (1) access disparities due to technical literacy and resource constraints Wilson et al. (2003); Sanders and Scanlon (2021); Lythreatis et al. (2022), as sophisticated personalization may primarily benefit tech-savvy users while marginalizing others; and (2) social polarization Weidinger et al. (2022); Shelby et al. (2022) through selective exposure and information asymmetry Abdelzaher et al. (2020); Gurkan et al. (2024), where personalized content delivery can create echo chambers and fragment public discourse.
To mitigate these risks, both technical and operational measures are crucial. Technical solutions include federated preference learning to protect user privacy, differential privacy mechanisms for preference data, and robust debiasing techniques during model training. Operational guidelines should encompass regular bias auditing, transparent preference management policies, and mechanisms for users to understand and control their preference profiles. However, implementing these safeguards while maintaining personalization effectiveness remains an ongoing challenge.